# A simplified proof of the CSP Dichotomy Conjecture and XY-symmetric operations

### v1

@@ 4.2 Proof of the main result
4.3 Proof of Theorems 44 and 45 (Finding a reduction)
@@ 4.4 Proof of Theorem 46 (Fixing an operation)
Let us show that $g$ satisfies the required properties (2)-(5). Since $f^{(1)}$ is XY-symmetric, by Lemmas 52 and 54, $g^{(1)}$ is also XY-symmetric. Property (3) follows from Lemma 55. Let us prove property (4). Consider two tuples $\delta,\delta^{\prime}\in\{c,d\}^{n}$ such that $\delta(1)=\delta^{\prime}(1)=c$, $N_{d}(\delta)=N_{d}(\delta^{\prime})$, and $(c,d)\in P$. We need to prove that $g^{(2,1)}(\delta)=g^{(2,1)}(\delta^{\prime})$. Since $(c,d)\in P$, by Lemma 54 for every $\gamma\in T_{a,b}^{n,k+1}$ we have $t_{0,\xi(\gamma,\alpha)}^{f,\xi(\gamma,\beta)}(\delta)=t_{0}^{f}(\delta)$ and $t_{0,\xi(\gamma,\alpha)}^{f,\xi(\gamma,\beta)}(\delta^{\prime})=t_{0}^{f}(\delta^{\prime})$. Hence $g^{(2,1)}(\delta)-g^{(2,1)}(\delta^{\prime})=f^{(2,1)}(\delta)-f^{(2,1)}(\delta^{\prime})=0$.
@@ 5.2 Subuniverses of types $\mathcal{BA},\mathcal{C},\mathcal{S}$
First, let us show that $C\neq\varnothing$. For every $n\geqslant 2$ put $W_{n}=\{(a_{1},\dots,a_{n})\mid\exists b\forall i\colon R(a_{i},b)\}$. If $W_{|A|}=A^{|A|}$, then $C\neq\varnothing$. Otherwise, choose the minimal $n\geqslant 2$ such that $W_{n}\neq A^{n}$. Since $\operatorname{LeftLinked}(R)=A^{2}$, $\operatorname{LeftLinked}(W_{n})=A^{2}$. Looking at $W_{n}$ as at binary relation $W_{n}\leqslant A\times A^{n-1}$ and using Lemmas 69 and 66 we derive the existence of BA or central subuniverse on $A$, which contradicts our assumptions.
@@ 5.3 Intersection property
Subsubcase 1B1. $T_{n}\in\{\mathcal{BA},\mathcal{C}\}$. Combining Lemmas 61 and 64, we obtain that $S_{n}\cap(B_{k-1}/\sigma_{k})^{|A|}<_{T_{n}}(B_{k-1}/\sigma_{k})^{|A|}$, which by Lemma 66 implies the existence of BA or central subuniverse on $B_{k-1}/\sigma_{k}$ and contradicts the definition of a dividing congruence.
Subsubcase 1B2. $T_{n}=\mathcal{S}$. Choose $G\leqslant B_{n}$ such that $G<_{\mathcal{BA},\mathcal{C}}B_{n-1}$. By Lemma 63 $G\cap B_{k}\neq\varnothing$, hence even if we replace $B_{n}$ by $G$ in the definition of $S_{n}$ the intersection $S_{n}\cap(B_{k-1}/\sigma_{k})^{|A|}$ will be nonempty. Then Lemmas 61 and 64 imply the existence of both BA and central subuniverse on $B_{k-1}/\sigma_{k}$ and contradicts the definition of a dividing congruence.
@@
Subcase 2A. $T_{k}\in\{\mathcal{BA},\mathcal{C}\}$. Combining Lemmas 61 and 64 we obtain $((B_{k}\circ\delta)\cap B_{m-1})/\sigma_{m}\leqslant_{T_{k}}((B_{k-1}\circ\delta)\cap B_{m-1})/\sigma_{m}$, which by the definition of a dividing congruence implies $((B_{k}\circ\delta)\cap B_{m-1})/\sigma_{m}=((B_{k-1}\circ\delta)\cap B_{m-1})/\sigma_{m}=B_{m-1}/\sigma_{m}$ and completes this case.
@@
Subsubcase 2A1. $T_{m}\in\{\mathcal{BA},\mathcal{C}\}$. Combining Lemmas 61 and 64, we obtain $S_{m,0}\cap(B/\delta)^{|A|}<_{T_{m}}(B/\delta)^{|A|}$, which by Lemma 66 implies the existence of BA or central subuniverse on $B/\delta$ and contradicts the definition of a dividing congruence.
Subsubcase 2A2. $T_{m}=\mathcal{S}$. Choose $G\leqslant B_{m}$ such that $G<_{\mathcal{BA},\mathcal{C}}B_{m-1}$. By Lemma 63 $G\cap B\neq\varnothing$, hence even if we replace $B_{m}$ by $G$ in the definition of $S_{m,0}$ the intersection $S_{m,0}\cap(B/\delta)^{|A|}$ will be nonempty. Then Lemmas 61 and 64 imply the existence of both BA and central subuniverse on $B/\delta$ and contradicts the definition of a dividing congruence.
@@
Subsubcase 2B1. $\mathcal{T}_{n}\in\{\mathcal{BA},\mathcal{C}\}$. Combining Lemmas 61 and 64, we obtain $S_{k,n}\cap(B/\delta)^{|A|}<_{\mathcal{T}_{n}}(B/\delta)^{|A|}$, which by Lemma 66 implies the existence of BA or central subuniverse on $B/\delta$ and contradicts the definition of a dividing congruence.
Subsubcase 2B2. $\mathcal{T}_{n}=\mathcal{S}$. Choose $G\leqslant C_{n}$ such that $G<_{\mathcal{BA},\mathcal{C}}C_{n-1}$. By the inductive assumption $G\cap B_{k}\neq\varnothing$. Then even if we replace $C_{n}$ by $G$ in the definition of $S_{m,n}$ we get a nonempty intersection with $(B/\delta)^{|A|}$. By Lemmas 61 and 64 we obtain both BA and central subuniverse on $(B/\delta)^{|A|}$, which by Lemma 66 implies the existence of BA and central subuniverse on $B/\delta$ and contradicts the definition of a dividing congruence.
@@ 5.4 Properties of PC or Linear congruences
By Lemmas 81 and 82, $\mathbf{A}/\sigma$ is affine and, therefore, polynomially equivalent to an $\mathbf{R}$-module $\mathbf{G}=(G,+,0,-,(r)_{r\in\mathbf{R}})$. Composing the bridge $\delta$ with itself we get a bridge $\delta_{0}\supseteq\delta$ which is an equivalence relation on $\mathbf{A}^{2}$. Then $\delta_{0}/\sigma$ is a congruence on the $\mathbf{R}$-module $\mathbf{G}^{2}$. Congruences in $\mathbf{R}$-modules come from submodules ($u$ and $v$ are congruent if $u-v$ is in a submodule). Since the diagonal of $(A/\sigma)^{2}$ is the block of the congruence $\delta_{0}/\sigma$, the corresponding submodule is $\{(x,y):x=y\}$. Hence $\delta_{0}/\sigma=\{((x_{1},y_{1}),(x_{2},y_{2})):x_{1}-x_{2}=y_{1}-y_{2}\}$. It remains to show that $\delta=\delta_{0}$. Since $\delta/\sigma$ is preserved by the Maltsev operation $x-y+z$ and $\widetilde{\delta}=\operatorname{pr}_{1,2}(\delta)=A^{2}$, we derive that $\operatorname{pr}_{1,2,3}(\delta)=A^{3}$. Since the last element of $\delta/\sigma$ is uniquely determined by the first three, we derive $\delta=\delta_{0}$. ∎
@@ 5.5 Types interaction
Assume that $T_{1}\in\{\mathcal{BA},\mathcal{C}\}$ and $T_{2}=\mathcal{D}$. By Lemma 77(d), $(B_{1}\cap B_{2})/\sigma_{2}=C_{2}/\sigma_{2}$ or $(B_{1}\cap B_{2})/\sigma_{2}=B_{2}/\sigma_{2}$. First case would imply that $B_{1}\cap B_{2}=B_{1}\cap C_{2}$, which contradicts $C_{1}\cap C_{2}=\varnothing$ and $C_{1}\cap B_{2}\neq\varnothing$. Thus, we assume that $(B_{1}\cap B_{2})/\sigma_{2}=B_{2}/\sigma_{2}$. Combining Lemmas 62 and 64 we obtain $(C_{1}\cap B_{2})/\sigma_{2}<_{T_{1}}(B_{1}\cap B_{2})/\sigma_{2}=B_{2}/\sigma_{2}$, which contradicts the definition of a dividing congruence.
@@ 5.6 Factorization of strong subalgebras
If $T\in\{\mathcal{BA},\mathcal{C}\}$ then by Lemmas 62 and 64 $S^{\prime\prime}<_{T}S^{\prime}=(B_{1}/\sigma)^{|A_{1}|}$ and by Lemma 66 there exists a BA or central subuniverse on $B_{1}/\sigma$, which contradicts our assumptions.
If $T=\mathcal{S}$ then choose $D_{2}\leqslant B_{2}^{\prime\prime}$ such that $D_{2}<_{\mathcal{BA},\mathcal{C}}B_{2}^{\prime}$. Combining Corollary 73 and Lemma 77(s) we obtain that $R\cap(B_{1}\times D_{2})\neq\varnothing.$ Hence, Lemmas 62 and 64 imply the existence of both BA and central subuniverse on $S^{\prime}=(B_{1}/\sigma)^{|A_{1}|}$ and by Lemma 66 there exists a BA and central subuniverse on $B_{1}/\sigma$, which contradicts our assumptions.
@@
By the choice of $B_{2}^{\prime}$ there exists $d\in B_{2}^{\prime}/\delta$ such that $(B_{1}/\sigma)^{|A_{1}|}\times\{d\}\subseteq S$ but $(B_{1}/\sigma)^{|A_{1}|}\times B_{2}^{\prime\prime}/\delta\neq S$. Since $\operatorname{pr}_{|A_{1}|+1}(S)/\delta=\operatorname{pr}_{2}(R\cap(B_{1}\times B_{2}^{\prime}))/\delta$, Corollary 78 implies that $\operatorname{pr}_{|A_{1}|+1}(S)/\delta=B_{2}^{\prime}/\delta$. Hence, $S$ is a central relation. Combining Lemmas 68 and 66 we get a contradiction with the assumption that $B_{1}/\sigma$ and $B_{2}^{\prime}/\delta$ are BA and center free. ∎
@@
For $n=2$ it follows from Lemmas 88 and 7. Put $B_{2}^{\prime}=B_{2}\cap C_{3}\cap\dots\cap C_{n}$ and $C_{2}^{\prime}=C_{2}\cap B_{2}^{\prime}$. By Lemma 19 $C_{2}^{\prime}\leqslant_{T_{2}(\sigma_{2})}B_{2}^{\prime}\lll A$. Also we have $C_{1}\cap C_{2}^{\prime}=\varnothing$, $C_{1}\cap B_{2}^{\prime}\neq\varnothing$, and $B_{1}\cap C_{2}^{\prime}\neq\varnothing$. Then Lemmas 88 and 7 imply that $T_{1}=T_{2}$, and if $T_{1}\in\{\mathcal{PC},\mathcal{L}\}$ then there is a bridge $\delta$ from $\sigma_{1}$ to $\sigma_{2}$ such that $\widetilde{\delta}=\sigma_{1}\circ\sigma_{2}$. If $T_{1}=\mathcal{PC}$ then $\sigma_{1}\circ\sigma_{2}=\sigma_{1}=\sigma_{2}$, as otherwise composing the bridge $\delta$ with itself we would get a nontrivial bridge from $\sigma_{i}$ to $\sigma_{i}$. Additionally, this implies that $n$ cannot be greater than 2 for $T_{1}=\mathcal{PC}$, as in this case the intersection $C_{1}\cap C_{2}$ must be empty.

### v2

@@ 4.2 Proof of the main result
4.3 Proof of Theorems 45 and 46 (Finding a reduction)
@@ 4.4 Proof of Theorem 47 (Fixing an operation)
Let us show that $g$ satisfies the required properties (2)-(5). Since $f^{(1)}$ is XY-symmetric, by Lemmas 53 and 55, $g^{(1)}$ is also XY-symmetric. Property (3) follows from Lemma 56. Let us prove property (4). Consider two tuples $\delta,\delta^{\prime}\in\{c,d\}^{n}$ such that $\delta(1)=\delta^{\prime}(1)=c$, $N_{d}(\delta)=N_{d}(\delta^{\prime})$, and $(c,d)\in P$. We need to prove that $g^{(2,1)}(\delta)=g^{(2,1)}(\delta^{\prime})$. Since $(c,d)\in P$, by Lemma 55 for every $\gamma\in T_{a,b}^{n,k+1}$ we have $t_{0,\xi(\gamma,\alpha)}^{f,\xi(\gamma,\beta)}(\delta)=t_{0}^{f}(\delta)$ and $t_{0,\xi(\gamma,\alpha)}^{f,\xi(\gamma,\beta)}(\delta^{\prime})=t_{0}^{f}(\delta^{\prime})$. Hence $g^{(2,1)}(\delta)-g^{(2,1)}(\delta^{\prime})=f^{(2,1)}(\delta)-f^{(2,1)}(\delta^{\prime})=0$.
@@ 5.2 Subuniverses of types $\mathcal{BA},\mathcal{C},\mathcal{S}$
First, let us show that $C\neq\varnothing$. For every $n\geqslant 2$ put $W_{n}=\{(a_{1},\dots,a_{n})\mid\exists b\forall i\colon R(a_{i},b)\}$. If $W_{|A|}=A^{|A|}$, then $C\neq\varnothing$. Otherwise, choose the minimal $n\geqslant 2$ such that $W_{n}\neq A^{n}$. Since $\operatorname{LeftLinked}(R)=A^{2}$, $\operatorname{LeftLinked}(W_{n})=A^{2}$. Looking at $W_{n}$ as at binary relation $W_{n}\leqslant A\times A^{n-1}$ and using Lemmas 70 and 67 we derive the existence of BA or central subuniverse on $A$, which contradicts our assumptions.
@@ 5.3 Intersection property
Subsubcase 1B1. $T_{n}\in\{\mathcal{BA},\mathcal{C}\}$. Combining Lemmas 62 and 65, we obtain that $S_{n}\cap(B_{k-1}/\sigma_{k})^{|A|}<_{T_{n}}(B_{k-1}/\sigma_{k})^{|A|}$, which by Lemma 67 implies the existence of BA or central subuniverse on $B_{k-1}/\sigma_{k}$ and contradicts the definition of a dividing congruence.
Subsubcase 1B2. $T_{n}=\mathcal{S}$. Choose $G\leqslant B_{n}$ such that $G<_{\mathcal{BA},\mathcal{C}}B_{n-1}$. By Lemma 64 $G\cap B_{k}\neq\varnothing$, hence even if we replace $B_{n}$ by $G$ in the definition of $S_{n}$ the intersection $S_{n}\cap(B_{k-1}/\sigma_{k})^{|A|}$ will be nonempty. Then Lemmas 62 and 65 imply the existence of both BA and central subuniverse on $B_{k-1}/\sigma_{k}$ and contradicts the definition of a dividing congruence.
@@
Subcase 2A. $T_{k}\in\{\mathcal{BA},\mathcal{C}\}$. Combining Lemmas 62 and 65 we obtain $((B_{k}\circ\delta)\cap B_{m-1})/\sigma_{m}\leqslant_{T_{k}}((B_{k-1}\circ\delta)\cap B_{m-1})/\sigma_{m}$, which by the definition of a dividing congruence implies $((B_{k}\circ\delta)\cap B_{m-1})/\sigma_{m}=((B_{k-1}\circ\delta)\cap B_{m-1})/\sigma_{m}=B_{m-1}/\sigma_{m}$ and completes this case.
@@
Subsubcase 2A1. $T_{m}\in\{\mathcal{BA},\mathcal{C}\}$. Combining Lemmas 62 and 65, we obtain $S_{m,0}\cap(B/\delta)^{|A|}<_{T_{m}}(B/\delta)^{|A|}$, which by Lemma 67 implies the existence of BA or central subuniverse on $B/\delta$ and contradicts the definition of a dividing congruence.
Subsubcase 2A2. $T_{m}=\mathcal{S}$. Choose $G\leqslant B_{m}$ such that $G<_{\mathcal{BA},\mathcal{C}}B_{m-1}$. By Lemma 64 $G\cap B\neq\varnothing$, hence even if we replace $B_{m}$ by $G$ in the definition of $S_{m,0}$ the intersection $S_{m,0}\cap(B/\delta)^{|A|}$ will be nonempty. Then Lemmas 62 and 65 imply the existence of both BA and central subuniverse on $B/\delta$ and contradicts the definition of a dividing congruence.
@@
Subsubcase 2B1. $\mathcal{T}_{n}\in\{\mathcal{BA},\mathcal{C}\}$. Combining Lemmas 62 and 65, we obtain $S_{k,n}\cap(B/\delta)^{|A|}<_{\mathcal{T}_{n}}(B/\delta)^{|A|}$, which by Lemma 67 implies the existence of BA or central subuniverse on $B/\delta$ and contradicts the definition of a dividing congruence.
Subsubcase 2B2. $\mathcal{T}_{n}=\mathcal{S}$. Choose $G\leqslant C_{n}$ such that $G<_{\mathcal{BA},\mathcal{C}}C_{n-1}$. By the inductive assumption $G\cap B_{k}\neq\varnothing$. Then even if we replace $C_{n}$ by $G$ in the definition of $S_{m,n}$ we get a nonempty intersection with $(B/\delta)^{|A|}$. By Lemmas 62 and 65 we obtain both BA and central subuniverse on $(B/\delta)^{|A|}$, which by Lemma 67 implies the existence of BA and central subuniverse on $B/\delta$ and contradicts the definition of a dividing congruence.
@@ 5.4 Properties of PC or Linear congruences
By Lemmas 82 and 83, $\mathbf{A}/\sigma$ is affine and, therefore, polynomially equivalent to an $\mathbf{R}$-module $\mathbf{G}=(G,+,0,-,(r)_{r\in\mathbf{R}})$. Composing the bridge $\delta$ with itself we get a bridge $\delta_{0}\supseteq\delta$ which is an equivalence relation on $\mathbf{A}^{2}$. Then $\delta_{0}/\sigma$ is a congruence on the $\mathbf{R}$-module $\mathbf{G}^{2}$. Congruences in $\mathbf{R}$-modules come from submodules ($u$ and $v$ are congruent if $u-v$ is in a submodule). Since the diagonal of $(A/\sigma)^{2}$ is the block of the congruence $\delta_{0}/\sigma$, the corresponding submodule is $\{(x,y):x=y\}$. Hence $\delta_{0}/\sigma=\{((x_{1},y_{1}),(x_{2},y_{2})):x_{1}-x_{2}=y_{1}-y_{2}\}$. It remains to show that $\delta=\delta_{0}$. Since $\delta/\sigma$ is preserved by the Maltsev operation $x-y+z$ and $\widetilde{\delta}=\operatorname{pr}_{1,2}(\delta)=A^{2}$, we derive that $\operatorname{pr}_{1,2,3}(\delta)=A^{3}$. Since the last element of $\delta/\sigma$ is uniquely determined by the first three, we derive $\delta=\delta_{0}$. ∎
@@ 5.5 Types interaction
Assume that $T_{1}\in\{\mathcal{BA},\mathcal{C}\}$ and $T_{2}=\mathcal{D}$. By Lemma 78(d), $(B_{1}\cap B_{2})/\sigma_{2}=C_{2}/\sigma_{2}$ or $(B_{1}\cap B_{2})/\sigma_{2}=B_{2}/\sigma_{2}$. First case would imply that $B_{1}\cap B_{2}=B_{1}\cap C_{2}$, which contradicts $C_{1}\cap C_{2}=\varnothing$ and $C_{1}\cap B_{2}\neq\varnothing$. Thus, we assume that $(B_{1}\cap B_{2})/\sigma_{2}=B_{2}/\sigma_{2}$. Combining Lemmas 63 and 65 we obtain $(C_{1}\cap B_{2})/\sigma_{2}<_{T_{1}}(B_{1}\cap B_{2})/\sigma_{2}=B_{2}/\sigma_{2}$, which contradicts the definition of a dividing congruence.
@@ 5.6 Factorization of strong subalgebras
If $T\in\{\mathcal{BA},\mathcal{C}\}$ then by Lemmas 63 and 65 $S^{\prime\prime}<_{T}S^{\prime}=(B_{1}/\sigma)^{|A_{1}|}$ and by Lemma 67 there exists a BA or central subuniverse on $B_{1}/\sigma$, which contradicts our assumptions.
If $T=\mathcal{S}$ then choose $D_{2}\leqslant B_{2}^{\prime\prime}$ such that $D_{2}<_{\mathcal{BA},\mathcal{C}}B_{2}^{\prime}$. Combining Corollary 74 and Lemma 78(s) we obtain that $R\cap(B_{1}\times D_{2})\neq\varnothing.$ Hence, Lemmas 63 and 65 imply the existence of both BA and central subuniverse on $S^{\prime}=(B_{1}/\sigma)^{|A_{1}|}$ and by Lemma 67 there exists a BA and central subuniverse on $B_{1}/\sigma$, which contradicts our assumptions.
@@
By the choice of $B_{2}^{\prime}$ there exists $d\in B_{2}^{\prime}/\delta$ such that $(B_{1}/\sigma)^{|A_{1}|}\times\{d\}\subseteq S$ but $(B_{1}/\sigma)^{|A_{1}|}\times B_{2}^{\prime\prime}/\delta\neq S$. Since $\operatorname{pr}_{|A_{1}|+1}(S)/\delta=\operatorname{pr}_{2}(R\cap(B_{1}\times B_{2}^{\prime}))/\delta$, Corollary 79 implies that $\operatorname{pr}_{|A_{1}|+1}(S)/\delta=B_{2}^{\prime}/\delta$. Hence, $S$ is a central relation. Combining Lemmas 69 and 67 we get a contradiction with the assumption that $B_{1}/\sigma$ and $B_{2}^{\prime}/\delta$ are BA and center free. ∎
@@
For $n=2$ it follows from Lemmas 89 and 8. Put $B_{2}^{\prime}=B_{2}\cap C_{3}\cap\dots\cap C_{n}$ and $C_{2}^{\prime}=C_{2}\cap B_{2}^{\prime}$. By Lemma 20 $C_{2}^{\prime}\leqslant_{T_{2}(\sigma_{2})}B_{2}^{\prime}\lll A$. Also we have $C_{1}\cap C_{2}^{\prime}=\varnothing$, $C_{1}\cap B_{2}^{\prime}\neq\varnothing$, and $B_{1}\cap C_{2}^{\prime}\neq\varnothing$. Then Lemmas 89 and 8 imply that $T_{1}=T_{2}$, and if $T_{1}\in\{\mathcal{PC},\mathcal{L}\}$ then there is a bridge $\delta$ from $\sigma_{1}$ to $\sigma_{2}$ such that $\widetilde{\delta}=\sigma_{1}\circ\sigma_{2}$. If $T_{1}=\mathcal{PC}$ then $\sigma_{1}\circ\sigma_{2}=\sigma_{1}=\sigma_{2}$, as otherwise composing the bridge $\delta$ with itself we would get a nontrivial bridge from $\sigma_{i}$ to $\sigma_{i}$. Additionally, this implies that $n$ cannot be greater than 2 for $T_{1}=\mathcal{PC}$, as in this case the intersection $C_{1}\cap C_{2}$ must be empty.